# Path planning with moving obstacles using stochastic optimal control

### v2

@@ Abstract
Navigating a collision-free and optimal trajectory for a robot is a challenging task, particularly in environments with moving obstacles such as humans. We formulate this problem as a stochastic optimal control problem. Since solving the full problem is computationally demanding, we introduce a tractable approximation whose Bellman equation can be solved efficiently. The resulting value function is then incorporated as a terminal penalty in an online rollout framework. We construct a trade-off curve between safety and performance to identify an appropriate weighting between them, and compare the performance with other methods. Simulation results show that our method reduces the time to target by a factor of two compared to the control barrier function method, while maintaining a larger minimum distance to the obstacle.
@@ I INTRODUCTION
Path planning is inherently a multi-objective task [4]. In particular, two common and often competing objectives are safety and performance [17]. In this paper, we introduce two cost terms: one associated with safety and the other with performance, and seek an appropriate trade-off between these objectives.
We will compare our method with the Control Barrier Function (CBF) approach to stochastic discrete time systems [16], and a receding horizon $A^{\star}$ algorithm. Numerical results show that the proposed method can outperform the CBF, and the receding horizon $A^{*}$ methods and achieves a balanced trade-off between the competing objectives. More precisely, our method reduces the expected time to reach the target by approximately a factor of two compared to CBF, while ensuring a larger expected minimum distance from the obstacle.
@@
    We formulate an optimal path planning problem as a stochastic optimal control problem, that outperforms other shortest path formulations on average.
@@ III Geometric Symmetry
where $F_{e}:\mathbb{R}_{+}\times\mathcal{U}\rightarrow\mathbb{R}_{+}$, $F_{\theta}:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]\times\mathcal{U}\times\mathcal{W}\rightarrow[0,\pi]$, and $F_{d}:\mathbb{R}_{+}\times[0,\pi]\times\mathcal{U}\times\mathcal{W}\rightarrow\mathbb{R}_{+}$, and $\bar{F}=\begin{bmatrix}F_{d}&F_{\theta}&F_{d}\end{bmatrix}^{T}$, and where
@@
V Effect of Probability Distribution and Constraints
[…]
where $\bar{w}_{l}=\mathbb{E}\left[w_{l}\right]$.
VI Numerical Results
@@
We perform $50$ independent simulations, where in each trial the target position is sampled uniformly at random from the grid. The initial positions of the robot and the moving obstacle are also sampled uniformly at random, subject to the constraints $r_{0}-t||_{2}>1$ and $h_{0}-r_{0}||_{2}>1$. For each set of initial conditions, we generate $10$ independent realizations of the stochastic disturbance $w$ to approximate the costs in (16) and (17), i.e., $V_{N,\lambda}^{\star}(h_{0},r_{0},t)$ and $V_{N\text{(CE)},\lambda}^{\star}(h_{0},r_{0},t)$. For the CBF method, we evaluate the performance for multiple values of $\alpha\in(0,1)$, and $d_{0}$. Fig. 3 shows the trade-off curve obtained by varying the weighting parameter $\lambda$ compared with the receding horizon $A^{*}$, CBF, and CBF (CE) methods. Recall that we maximize the expected minimum distance between the robot and the obstacle. To obtain a standard Pareto trade-off plot, where both axes represent minimization, we display its negative value in Fig. 3. Thus, points closer to the origin correspond to better performance in both objectives. From Fig. 3, we observe that the best trade-off between the competing costs occurs at $\lambda^{*}=5\times 10^{-6}$. Additionally, from this figure, the proposed method achieves a lower trade-off curve on average over the grid, compared to CBF and CBF (CE). This means that, for different horizon lengths, the rollout solutions lie in the lower-left region of the plot relative to the curves of CBF and CBF (CE), indicating better performance for both objectives. Moreover, Fig. 3 shows that increasing the horizon shifts the trade-off curve further down and to the left, meaning that the longer horizon yields a better trade-off between the objectives (lower values for both costs) compared to curves that lie above and to the right. Furthermore, from this figure, the receding horizon $A^{*}$ approach always yields a solution that prioritizes minimizing the expected time to reach the target, as its points lie toward the lower-right part of the plot.
@@
Figure 3: Trade-off curve between expected time to target (horizontal axis) and negative minimum distance to the obstacle (vertical axis) for different horizons, compared with receding horizon $A^{*}$, CBF, and CBF (CE) over the grid
Table I summarizes the expected time to reach the target and the expected minimum distance to the target for favorable values of $\gamma$, $\alpha$, and $d_{0}$ obtained from Fig. 3. From this table, the receding horizon $A^{*}$ method provides the fastest time to reach the target, while also achieving the smallest minimum distance among all methods. In contrast, the CBF and CBF (CE) methods exhibit a larger expected time to reach the target. Using the proposed method with different horizon lengths results in a balance between reaching the target efficiently and maintaining a larger minimum expected distance to the target, illustrating the trade-off between performance and safety.
@@
Now, we let $h_{0}=\begin{bmatrix}2&6\end{bmatrix}^{T}$, $r_{0}=\begin{bmatrix}4&12\end{bmatrix}^{T}$, and $t=\begin{bmatrix}4&3\end{bmatrix}^{T}$. We consider $100$ different realizations of $w$, and compare the proposed method with the receding-horizon $A^{\star}$, CBF, and CBF (CE) approaches. Fig. 4 illustrates the trade-off curve obtained by varying the weighting parameter $\lambda$ compared with the other methods for the given initial condition and for different combinations of $\alpha\in(0,1)$, and $d_{0}$. From Fig. 4, the best trade-off between the competing costs occurs at $\lambda^{*}=5\times 10^{-6}$, which is the same value that we obtained from Fig. 3. Note that Fig. 3 shows the trade-off curve computed over the entire grid, whereas Fig. 4 illustrates the trade-off curve corresponding to the given initial conditions. From Fig. 4, for the given initial condition, the proposed method achieves a lower trade-off curve than CBF and CBF (CE) for some values of $\lambda$, while for other values of $\lambda$ the curve lies above those of CBF and CBF (CE). This indicates that, the proposed method does not always outperforms the other methods. However, when averaged over the entire grid of initial conditions, as shown in Fig. 3, our method outperforms the receding horizon $A^{*}$, CBF, and CBF (CE) methods.
@@
Figure 4: Trade-off curve between expected time to target (horizontal axis) and negative minimum distance to the obstacle (vertical axis) for different horizons, for different horizons compared with receding-horizon $A^{*}$, CBF, and CBF (CE) for the given initial conditions
@@
From Fig. 4, we select $\lambda=\lambda^{\ast}=5\times 10^{-6}$ for our method, and $\alpha=0.75$ and $d_{0}=1$ for the CBF and CBF (CE) methods, as these values provide the best trade-off for the given initial condition. We then plot the trajectory and control signal for a sample run of the different methods. Fig. 5 shows the trajectory of the robot and the moving obstacle for the different methods. From this figure, the trajectory of the robot when using the proposed approach makes a detour to avoid the moving obstacle. The trajectories produced by the CBF and CBF (CE) methods are identical in this example, and they also take a detour in order to satisfy the barrier function constraint. In contrast, the receding-horizon $A^{*}$ approach does not account for the moving obstacle and moves directly toward the goal without considering it. Fig. 6 shows the distance between the robot and the target, and the distance between the robot and the moving obstacle, over time for the different methods. From this figure, we notice that the robot reaches the target faster under the $A^{*}$ approach than under our proposed method, CBF, or CBF (CE). However, the distance between the robot and the moving obstacle is generally larger when using our proposed method compared to the $A^{*}$ approach, and is comparable to the distances achieved by the CBF and CBF (CE) methods. These results indicate that our method performs better than the other methods.
@@ VII Conclusions
In this paper, we used geometric symmetry to solve path planning problem for a robot in an environment with a stochastically moving obstacle. Compared with receding-horizon $A^{*}$, CBF, and CBF (CE), the proposed method achieves a better trade-off between reaching the target and maintaining safety; specifically, the trade-off curves lie below those of the other methods. Future work includes extending this framework to three-dimensional environments or to scenarios with moving targets. In the three-dimensional case with a static target, the symmetry reduction decreases the state dimension from $6$ to $3$. Additional directions include applying the reduced dimension value function within a reinforcement learning framework to learn the optimal value function in unknown environments, and generalizing the approach to multiple moving obstacles.
TABLE I: Comparison of proposed method with receding horizon $A^{\star}$, CBF, and CBF (CE), for $\lambda=5\times 10^{-6}$, $\alpha=0.75$, $d_{0}=1$
Method Expected Time to Reach the Target Expected Minimum Distance to Obstacle
Rollout ($N=1$) $10.278$ $8.0034$
Rollout ($N=1$, CE) $10.67$ $8.0377$
Rollout ($N=2$, CE) $11.388$ $8.2199$
Rollout ($N=3$, CE) $10.804$ $8.2337$
CBF $29.824$ $7.2857$
CBF (CE) $29.14$ $7.3081$
Receding Horizon $A^{\star}$ $8.896$ $7.3332$
@@ VII-A Symmetry of Optimal Value Function in Stochastic Shortest Path Problems
    We assume that there exists at least one proper policy from the initial states—i.e., a policy that reaches $t$ with probability one and finite expected cost. Under this assumption, the Bellman operator admits a unique minimal nonnegative fixed point, [21, Proposition 1]:
@@
Substituting $y$ with $x$ in (36) and combining with (32) yields the desired invariance (20). ∎
@@ VII-B Dimension Reduction
For all $\alpha\in G$ we have $\rho(\phi(\alpha,x))=\rho(x)$[23, Lemma 1], meaning that $\rho$ is invariant under the group action. Furthermore, the restriction of $\rho$ to $\mathcal{C}$ is injective. For additional details on this method, we refer the reader to [11, 23].
@@ VII-C Dynamics of Reduced-States
where $F_{\theta}:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]\times\mathcal{U}\times\mathcal{W}\rightarrow[0:\pi]$, and where

### v3

@@ Abstract
Navigating a collision-free and optimal trajectory for a robot is a challenging task, particularly in environments with moving obstacles such as humans. We formulate this problem as a stochastic optimal control problem. Since solving the full problem is computationally demanding, we introduce a tractable approximation whose Bellman equation can be solved efficiently. The resulting value function is then incorporated as a terminal penalty in an online rollout framework. We construct a trade-off curve between safety and performance to identify an appropriate weighting between them, and compare the performance with other methods. Simulation results show that the proposed rollout approach can be tuned to reach the target in nearly the same expected time as receding horizon $A^{\star}$ while maintaining a larger expected minimum distance to the moving obstacle. The results also show that the proposed method outperforms the considered CBF-based methods when a larger obstacle clearance is desired, while achieving comparable performance otherwise.
@@ I Introduction
Path planning is inherently a multi-objective task [4]. In particular, two common and often competing objectives are safety and performance [17]. In this paper, we introduce two cost terms: one associated with safety and the other with performance, and seek an appropriate trade-off between these objectives. An important feature of the proposed formulation is that this trade-off can be tuned explicitly through a single parameter $\lambda$, which gives a simple mechanism for selecting a policy with a desired balance between performance and safety.
We compare our method with a stochastic discrete time CBF approach [16] and a receding horizon $A^{\star}$ method. A key advantage of the proposed method is that the parameter $\lambda$ provides a direct and interpretable way to tune the trade-off between time to target and obstacle clearance. This makes it straightforward to generate a trade-off curve and to select a desired operating point. In contrast, for CBF-based methods, performance depends strongly on the choice of barrier function and associated tuning parameters, and obtaining a desired trade-off is less direct. Our numerical results show that the proposed method can achieve nearly the same expected time to target as receding horizon $A^{\star}$ while maintaining a larger expected minimum distance to the obstacle. They also show that the proposed method outperforms the considered CBF-based methods when a larger obstacle clearance is desired, while providing comparable performance otherwise.
@@
    We formulate an optimal path planning problem as a stochastic optimal control problem and develop an efficient approximate solution suitable for online rollout.
@@
  3. 3.
    We show through simulations that the proposed method provides an easily tunable trade-off between time to target and obstacle clearance, while comparing favorably with receding horizon $A^{\star}$ and the considered CBF-based methods.
@@ III Geometric Symmetry
where $F_{e}:\mathbb{R}_{+}\times\mathcal{U}\rightarrow\mathbb{R}_{+}$, $F_{\theta}:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]\times\mathcal{U}\times\mathcal{W}\rightarrow[0,\pi]$, and $F_{d}:\mathbb{R}_{+}\times[0,\pi]\times\mathcal{U}\times\mathcal{W}\rightarrow\mathbb{R}_{+}$, and $\bar{F}=\begin{bmatrix}F_{d}&F_{e}&F_{\theta}\end{bmatrix}^{T}$, and where
@@
V Effect of Probability Distribution and Constraints
[…]
where $\bar{w}_{l}=\mathbb{E}\left[w_{l}\right]$.
VI Numerical Results
@@
We perform $50$ independent simulations, where in each trial the target position is sampled uniformly at random from the grid. The initial positions of the robot and the moving obstacle are also sampled uniformly at random, subject to the constraints $r_{0}-t||_{2}>1$ and $h_{0}-r_{0}||_{2}>1$. For each set of initial conditions, we generate $10$ independent realizations of the stochastic disturbance $w$ to approximate the costs in (16) and (17), i.e., $V_{N,\lambda}^{\star}(h_{0},r_{0},t)$ and $V_{N\text{(CE)},\lambda}^{\star}(h_{0},r_{0},t)$. For the CBF method, we evaluate the performance for multiple values of $\alpha\in(0,1)$, and $d_{0}$. Fig. 3 shows the trade-off curve obtained by varying the weighting parameter $\lambda$ compared with the receding horizon $A^{\star}$, CBF, and CBF (CE) methods. Recall that we maximize the expected minimum distance between the robot and the obstacle. To obtain a standard Pareto trade-off plot, where both axes represent minimization, we display its negative value in Fig. 3. Thus, points closer to the origin correspond to better performance in both objectives. From Fig. 3, the proposed method provides a favorable trade-off between expected time to target and expected minimum obstacle distance over the grid of initial conditions. In particular, for operating points with larger desired obstacle clearance, the proposed method achieves better trade-offs than the considered CBF and CBF (CE) methods. For smaller desired clearance, the performance is comparable. Moreover, the figure shows that the proposed method can be tuned to operate close to the $A^{\star}$ solution in expected time to target while maintaining a larger expected minimum distance to the obstacle. It is possible that a different barrier function and tuning could improve the CBF trade-off. However, identifying such a barrier function is not straightforward. In contrast, the proposed method offers a simple and direct tuning mechanism through $\lambda$, which makes it easy to explore and select the desired time-safety trade-off.
Moreover, Fig. 3 shows that increasing the horizon shifts the trade-off curve further down and to the left, meaning that the longer horizon yields a better trade-off between the objectives (lower values for both costs) compared to curves that lie above and to the right. Furthermore, from this figure, the receding horizon $A^{\star}$ approach yields a solution that prioritizes minimizing the expected time to reach the target, as its points lie toward the lower-right part of the plot.
@@
Figure 3: Trade-off curve between expected time to target (horizontal axis) and negative minimum distance to the obstacle (vertical axis) for different horizons, compared with receding horizon $A^{\star}$, CBF, and CBF (CE) over the grid
@@
Now, we let $h_{0}=\begin{bmatrix}2&6\end{bmatrix}^{T}$, $r_{0}=\begin{bmatrix}4&12\end{bmatrix}^{T}$, and $t=\begin{bmatrix}4&3\end{bmatrix}^{T}$. We consider $100$ different realizations of $w$, and compare the proposed method with the receding-horizon $A^{\star}$, CBF, and CBF (CE) approaches. Fig. 4 illustrates the trade-off curve obtained by varying the weighting parameter $\lambda$ compared with the other methods for the given initial condition and for different combinations of $\alpha\in(0,1)$, and $d_{0}$. From Fig. 4, the best trade-off between the competing costs occurs at $\lambda^{*}=5\times 10^{-6}$, which is the same value that we obtained from Fig. 3. Note that Fig. 3 shows the trade-off curve computed over the entire grid, whereas Fig. 4 illustrates the trade-off curve corresponding to the given initial conditions. From Fig. 4, for the given initial condition, the proposed method outperforms CBF and CBF (CE) in the region where a larger minimum obstacle distance is desired, while achieving comparable performance in the rest of the trade-off range. Thus, the proposed method is particularly advantageous when safety, in the sense of larger obstacle clearance, is prioritized.
@@
Figure 4: Trade-off curve between expected time to target (horizontal axis) and negative minimum distance to the obstacle (vertical axis) for different horizons, compared with receding-horizon $A^{\star}$, CBF, and CBF (CE) for the given initial conditions
@@
From Fig. 4, we select $\lambda=\lambda^{\ast}=5\times 10^{-6}$ for our method, and $\alpha=0.75$ and $d_{0}=1$ for the CBF and CBF (CE) methods, as these values provide the best trade-off for the given initial condition. We then plot the trajectory and control signal for a sample run of the different methods. Fig. 5 shows the trajectory of the robot and the moving obstacle for the different methods. From this figure, the trajectory of the robot when using the proposed approach makes a detour to avoid the moving obstacle. The trajectories produced by the CBF and CBF (CE) methods are identical in this example, and they also take a detour in order to satisfy the barrier function constraint. In contrast, the receding-horizon $A^{\star}$ approach does not account for the moving obstacle and moves directly toward the goal without considering it. Fig. 6 shows the distance between the robot and the target, and the distance between the robot and the moving obstacle, over time for the different methods. This figure shows that receding-horizon $A^{\star}$ reaches the target slightly faster than the proposed method. However, the proposed method maintains a larger distance from the moving obstacle during the motion. Hence, the proposed method can achieve time to target performance close to $A^{\star}$ while offering improved safety. Relative to CBF and CBF (CE), the proposed method achieves comparable performance in this example, while offering the practical advantage that its time-safety trade-off can be tuned directly through $\lambda$.
@@ VII Conclusions
In this paper, we proposed a symmetry reduced stochastic optimal control approach for path planning in the presence of a stochastically moving obstacle. Compared with receding horizon $A^{\star}$, CBF, and CBF (CE), the proposed method provides a simple and interpretable tuning mechanism through $\lambda$ to select and visualize the trade-off between time to target and obstacle clearance. The numerical results show that the proposed method can achieve nearly the same expected time to target as receding horizon $A^{\star}$ while maintaining a larger expected minimum distance to the obstacle. They also show that the proposed method outperforms the considered CBF based methods when a larger obstacle clearance is desired, while offering comparable performance otherwise. Future work includes extending this framework to three-dimensional environments or to scenarios with moving targets. In the three-dimensional case with a static target, the symmetry reduction decreases the state dimension from $6$ to $3$. Additional directions include applying the reduced dimension value function within a reinforcement learning framework to learn the optimal value function in unknown environments, and generalizing the approach to multiple moving obstacles.
@@ VII-A Symmetry of Optimal Value Function in Stochastic Shortest Path Problems
    We assume that there exists at least one proper policy from the initial states—i.e., a policy that reaches $t$ with probability one and finite expected cost. Under this assumption, the Bellman operator admits a unique minimal nonnegative fixed point, [20, Proposition 1]:
@@
Substituting $y$ with $x$ in (36) and combining with (32) yields the desired invariance (20).
@@ VII-B Dimension Reduction
For all $\alpha\in G$ we have $\rho(\phi(\alpha,x))=\rho(x)$[22, Lemma 1], meaning that $\rho$ is invariant under the group action. Furthermore, the restriction of $\rho$ to $\mathcal{C}$ is injective. For additional details on this method, we refer the reader to [11, 22].
@@ VII-C Dynamics of Reduced-States
where $F_{\theta}:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]\times\mathcal{U}\times\mathcal{W}\rightarrow[0,\pi]$, and where
@@
Acknowledgment
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.